OpenStudy (anonymous):

what are the x-intercepts of the graph of the equation y-x^2= -4x -5

OpenStudy (anonymous):

The x-intercepts occur where y=0 on a graph. This means if you put 0 in for y and solve the equation for x, you'll get your x-intercepts.

OpenStudy (anonymous):

0-x^2=-4x-5 -x^2=-4x-5 0=x^2-4x-5 Now factor the quadratic and set each piece equal to 0 to find your x values.